5A A Bernoulli Trials experiment consists of 4 trials, with a 4/5 probability of success on each trial.

What is the probability of at least 1 success and at least 1 failure?

What is the probability of 2 successes, given at least 1 success? What is the probability of at least 2 successes, given at least 2 failures?

Enter your answers as whole numbers or fractions in lowest terms.
Answer #1

P(X) = nCx px qn-x

n = 4

p = 4/5

q = 1/5

1. P(at least 1 success and at least 1 failure) = 1 - P(all successes) - P(all failures)

= 1 - (4/5)4 - (1/5)4

= 368/625

2. P(2 successes | at least 1 success) = P(2 successes) / P(at least one success)

= [4C2 x (4/5)2 x (1/5)2] / [1 - (1/5)4]

= 2/13

3. P(at least 2 successes | at least 2 failures) = P(2 success and 2 failures) / P(at least 2 failures)

= [4C2 x (4/5)2 x (1/5)2] / [P(0 success) + P(1 success) + P(2 success)]

= [4C2 x (4/5)2 x (1/5)2]/[(1/5)4 + 4 x (4/5)x(1/5)3 + 4C2 x (4/5)2 x (1/5)2]

= 96/113